OpenStudy (bruno102):

What is the equation of the ellipse with vertices (6, 0), (-6, 0) and co-vertices (0, 2), (0, -2)?

OpenStudy (bruno102):

Possible answers

OpenStudy (loser66):

how far you go?

OpenStudy (bruno102):

what do you mean?

OpenStudy (loser66):

I want to know where you got stuck.

OpenStudy (bruno102):

I have graphed each one, and non of them work for the question.

OpenStudy (bruno102):

At least from what I am doing.

OpenStudy (bruno102):

none*

OpenStudy (loser66):

the vertice (6,0) and (-6,0), covertices (0,2) and (0,-2), its graph should be

OpenStudy (loser66):



OpenStudy (bruno102):

Yes

OpenStudy (loser66):

and the standard form of this graph is??

OpenStudy (bruno102):

You mean the formula?

OpenStudy (loser66):

yup

OpenStudy (bruno102):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] ???

OpenStudy (loser66):

oh yea, I don't see the right answer from the options, either. hehehe...

OpenStudy (bruno102):

Really. That's what's on my assignment :(

OpenStudy (loser66):

The second one should have + in the middle , not - , then it turns true. I am sorry. It may be typo there. If it is not, then, I failed this question.

OpenStudy (bruno102):

That's is what I thought too. It's okay, I will have to ask my teacher haha. Thanks

OpenStudy (loser66):

Let tag some experts. @ganeshie8 @mathmale

OpenStudy (bruno102):

Okay :)

OpenStudy (loser66):

I don't give up that easy. hihihii

OpenStudy (bruno102):

Haha! :)

OpenStudy (loser66):

hihihi... I am kind of annoying Loser

OpenStudy (bruno102):

Noo

OpenStudy (loser66):

@ganeshie8 say something , please. Don't give me a medal for nothing

ganeshie8 (ganeshie8):

I think none of the options are right

OpenStudy (loser66):

lalala...thank you

OpenStudy (bruno102):

Yeah, that's what we thought.

ganeshie8 (ganeshie8):